# Unveiling Inefficiencies in LLM-Generated Code: Toward a Comprehensive Taxonomy

### v2

@@ Abstract
Large Language Models (LLMs) are widely adopted for automated code generation with promising results. Although prior research has assessed LLM-generated code and identified various quality issues- such as redundancy, poor maintainability, and sub-optimal performance- a systematic understanding and categorization of these inefficiencies remain unexplored. Therefore, we empirically investigate inefficiencies in LLM-generated code by state-of-the-art models, i.e., CodeLlama, DeepSeek-Coder, and CodeGemma. To do so, we manually analyze 492 generated code snippets in the HumanEval+ dataset. We then construct a taxonomy of inefficiencies in LLM-generated code that includes 5 categories (General Logic, Performance, Readability, Maintainability, and Errors) and 19 subcategories of inefficiencies. We validate the obtained taxonomy through an online survey with 58 LLM practitioners and researchers. The surveyed participants affirmed the completeness of the proposed taxonomy, and the relevance and the popularity of the identified code inefficiency patterns. Our qualitative findings indicate that inefficiencies are diverse and interconnected, affecting multiple aspects of code quality, with logic and performance-related inefficiencies being the most frequent and often co-occur while impacting overall code quality. Our taxonomy provides a structured basis for evaluating the quality of LLM-generated code and guiding future research to improve code generation efficiency.
@@
LLMs, Code Generation, Code Quality
@@ I Introduction
We conduct our study using HumanEval+  a widely used benchmark for code generation composed of 164 human-crafted Python tasks [6]. We selected three leading open-source LLM models -CodeLlama, DeepSeek-Coder, and CodeGemma- to generate code for the different tasks, resulting in 492 generated code snippets. To evaluate the quality and requirements conformity of the LLM-generated code, we employed the higher-capacity GPT-4o-mini model as a judge. The judgments provided initial guidelines for our manual exploratory and qualitative analysis to identify inefficiency patterns in LLM-generated code. We adopted open coding [18] to iteratively derive categories of inefficiencies. Our study results in a taxonomy spanning five categories: General Logic, Performance, Readability, Maintainability, and Errors, comprising 19 subcategories. We validate this taxonomy through a survey with 58 software practitioners and researchers who use LLMs for coding assistance.
@@
  • We study inefficiencies observed in LLM-generated code and their prevalence.
  • We propose a taxonomy of inefficiencies in LLM-generated code based on their characteristics.
@@ III Methodology
This section presents the methodology used in this study to construct and validate the taxonomy of inefficiencies in LLM-generated code. The process consists of five steps, as illustrated in Figure 1. First, we generated code using the HumanEval+ benchmark [6] and collected data from three open-source LLM4Code models: CodeLlama [4], DeepSeek-Coder [21], and CodeGemma [23]. Second, we leveraged a higher-capacity model (GPT-4o-mini) to judge the generated code and identify existing inefficiencies. In the third and fourth steps, we manually analyzed the code samples and the model’s judgments to develop a comprehensive taxonomy. Finally, in the fifth step, we conducted a survey with practitioners and researchers to validate the LLM-generated code inefficiencies identified in our taxonomy. The details of each step are presented in the remainder of this section.
@@ III-A Step 1: Data Selection and Collection
Regarding the data collection, considering the popularity and wide adoption of Python as a programming language [24], this study focused on Python-generated LLM code. Given the focus on Python, we selected HumanEval+ [6] (denoted as $\mathit{A}$ in Figure 1), a widely recognized and extensively employed benchmark for code generation [25, 26, 27, 28]. HumanEval+ enhances the original HumanEval [29] dataset by augmenting it with additional test cases, enabling a more rigorous and reliable assessment of the functional correctness of the generated code [6]. HumanEval+ comprises 146 Python programming tasks, each accompanied by (i) a task description provided as a doc string and (ii) associated unit tests to rigorously evaluate the correctness of the code snippets. Regarding the LLMs, we aimed to select those trained or fine-tuned for code generation tasks to enable a more meaningful and targeted analysis of inefficiencies. To ensure accessibility and feasibility, we established two selection criteria: (i) the models must be open source and (ii) have fewer than 8B parameters. Based on these criteria, we selected the three highest-performing models according to the HumanEval+ leaderboard [30]: CodeLlama-7B [31], GodeGemma-7B [32], and DeepSeek Coder-6.7B [33]. These models, beyond ranking among the top-performing models [30], are also widely adopted for code generation tasks [27]. Although our selection focuses on small open-source models, these models demonstrate performance comparable to proprietary ones, surpassing GPT-3.5 [21] and approaching GPT-4 [34], ensuring representativeness in our study. After defining the setup for our study, we proceeded with generating the code snippets. For the hyperparameters, we adopted a conservative approach by setting the temperature to 00, ensuring deterministic outputs and minimizing randomness in code generation. For prompt construction, we followed the default configuration recommended in the official Hugging Face documentation [32, 31, 33]. This approach ensured a fair assessment of inherent inefficiencies in LLM-generated code rather than artifacts caused by inadequate prompts. As a result of this step, we have a dataset of $164$ generated code snippets for each LLM, resulting in $492$ code snippets ($\mathit{B}$), from the initial dataset of tasks ($\mathit{A}$).
@@ IV-A2 Analysis of Inefficiency Patterns
Sub-Optimal Solution (Memory) 26.83 17.07 12.2 18.7 3.07 3.12
@@
3-Readability 6.71 6.1 1.22 4.67 3.22
Confusing Variable Naming 1.22 1.22 0.61 1.02 2.17 3
@@
Syntax Error 0 2.44 2.44 1.63 2.29 3.79
Missing Variable Declaration 0.61 0 0.61 0.41 2.17 3.60
Missing Module Import 6.1 3.66 1.22 3.66 2.86 3.26
@@
Readability (4.67%) and Errors (5.69%) were less frequent but non-trivial. Syntax Errors were rare ($<2.44\%$), confirming the strong syntactic grasp of code LLMs. However, Missing Imports (6.1% in CodeLlama) reveal persistent gaps in dependency resolution, a known limitation even in highly-efficient models like GPT-4 [53].
@@
To gain deeper insights into inefficiency relationships, we analyze how different categories co-occur across models. Figure 3 presents the normalized category co-occurrence heatmap across all models. The co-occurrence values are normalized row-wise by dividing each occurrence by the frequency of the corresponding category. Our results reveal strong inter-dependencies between inefficiency patterns in LLM-generated code. General Logic inefficiencies exhibit the highest co-occurrence rates with other categories, particularly Readability (0.74), Errors (0.64), and Maintainability (0.58). This suggests that flawed logic not only impacts correctness but also contributes to inefficiencies in clarity, structure, and long-term maintainability of the generated code. Similarly, Performance inefficiencies frequently co-occur with Readability (0.48) and Maintainability (0.46), highlighting that performance inefficiencies are often tied to code complexity and structure, rather than isolated algorithmic inefficiencies. Interestingly, Errors demonstrate weaker co-occurrences with other inefficiencies, with their strongest link being to Maintainability (0.36), indicating that while syntax and dependency errors are less common, they may still hinder maintainability efforts by requiring additional debugging and refactoring. Meanwhile, Readability and Maintainability are closely related (0.35), reinforcing the idea that clear code structures contribute to long-term maintainability. These findings suggest that improving LLM-generated logic and performance could lead to broader improvements across multiple inefficiency categories, reducing the need for manual corrections and enhancing code usability in real-world development.
@@ IV-B RQ2: Relevance of Identified Inefficiencies for Practitioners and Researchers
The survey remained open for four weeks, reaching $58$ responses, resulting in a response rate of $7.39$%, which is in line with rates observed in similar surveys in Software Engineering [39, 54, 55, 56]. First, we present the demographic information about the participants and their experience with LLM-assisted coding. Then, we provide a detailed analysis of the popularity and perceived relevance of inefficiencies in LLM-generated code reported by the participants.
@@ IV-B1 Demographic Information and Experience with Coding using LLMs
All participants answered the demographic questions. We had participants from both academia (60%) and industry (40%). The academic participants consisted of graduate students (29), undergraduate students (3), and postdoctoral researchers/lecturers (3). Industry participants included 11 software developers, 7 ML/data science developers, 4 senior executives, and 1 cybersecurity developer. About coding experience, 37 participants had more than 5 years of experience, while 11 had between 3-5 years, and 10 had between 1-3 years. LLM-assisted coding is widely adopted, with 98% of the participants using LLMs for coding. Among them, 68% rely exclusively on proprietary models, with ChatGPT and Copilot being the most popular (89.66% and 29.31%, respectively). Regarding open-source, the most popular models are Mistral, the Phi model family, and Qwen-Coder (5.17%, 3.45%, and 3.45%, respectively). Participants reported using a variety of programming languages, with the top five being Python, C/C++, JavaScript, Java, and Go (91.38%, 18.97%, 18.97%, 8.62%, and 3.45%, respectively).
@@
These inefficiencies focus on the challenges/difficulties that users face when interacting with LLMs, rather than the generated code itself. 6 participants reported that LLMs struggle with debugging or modifying specific parts of code, a finding consistent with previous studies [57]. In the same way, programming language and library popularity also play a significant role in the efficiency of LLM-generated code. 5 participants noted that LLMs generate more efficient code for popular languages and libraries like Python and NumPy compared to less common ones like DHL and Tinkter. This difference is likely because LLMs are trained on more examples of popular languages and libraries, making them more familiar with their syntax and usage. Programming language paradigms also play a role—P12111We refer to survey participants using the notation Px, where x is an assigned participant ID. observed that LLMs handle interpreted languages better than compiled ones. Finally, participants expressed varying preferences for chat-based LLMs (e.g., ChatGPT) versus inline code assistants (e.g., Copilot). While P41 found Copilot suggestions as “stupid suggestions and distracted”, P11 preferred inline assistance over chat-based interactions.
@@
Participants also reported various inefficiencies in the generated code that align with the focus of our taxonomy. 19 participants stated that LLMs often struggle to produce the correct logic on the first attempt, even for simple tasks with detailed descriptions, requiring multiple further iterations. P22 specifically noted that LLMs may generate “unnecessarily long or complex code (for simple tasks)” However, P15 found that for more complex tasks such as building a Flask application, LLMs tend to perform “oversimplification to the point of uselessness”, where generation is limited to high-level steps or generalized code that lack specificity, performance efficiency, and functionality. Additionally, 3 participants reported that handling exceptions and corner cases is often overlooked. These reported inefficiencies fall under the General Logic category of our taxonomy.
@@
Additionally, 6 participants highlighted performance issues. Others pointed to a lack of robust, modular, object-oriented architecture and insufficient documentation and comments as well. The inefficiencies reported by the participants align with the literature [16, 14, 16] and fall under Readability, Performance, and Maintainability categories of our taxonomy.
Besides inefficiencies, 7 participants identified context limitations as a key issue. 2 participants found that LLMs had a limited context window and frequently forgot past interactions, as mentioned in related work [47]. Meanwhile, P5 reported that LLMs tended to retain outdated context, stating that they needed to “start a new chat to get rid of hallucinations.”
@@ VI Threats to Validity
Internal Validity. A key internal threat to validity is the potential bias in the manual labeling of LLM-generated code. To mitigate such a threat, the first two authors independently labeled the generated code snippets. The inter-rater agreement at the category level is 0.846, which aligns with similar studies in software engineering [39, 13]. Additionally, to further validate our findings, we surveyed 58 participants from both academia and industry, ensuring a broader perspective on the identified inefficiencies. Additionally, we leveraged an LaaJ to assess generated code snippets, providing an additional layer of validation and reducing potential human bias in the labeling process. However, we recognize that LaaJ judgments may have missed certain inefficiencies or introduced biases that could influence human labeling. To mitigate this risk, the two labelers independently reviewed, refined, and expanded upon the LLM’s judgments, ensuring that inefficiencies were identified beyond those initially flagged by the model.

### v3

# A Taxonomy of Inefficiencies in LLM-Generated Python Code

@@ Abstract
Large Language Models (LLMs) are widely adopted for automated code generation with promising results. Although prior research has assessed LLM-generated code and identified various quality issues- such as redundancy, poor maintainability, and sub-optimal performance- a systematic understanding and categorization of these inefficiencies remain unexplored. Therefore, we empirically investigate inefficiencies in LLM-generated Python code by state-of-the-art models, i.e., CodeLlama, DeepSeek-Coder, and CodeGemma. To do so, we manually analyze 492 generated Python code snippets in the HumanEval+ dataset. We then construct a taxonomy of inefficiencies in LLM-generated Python code that includes 5 categories (General Logic, Performance, Readability, Maintainability, and Errors) and 19 subcategories of inefficiencies. We validate the obtained taxonomy through an online survey with 58 LLM practitioners and researchers. The surveyed participants affirmed the completeness of the proposed taxonomy, and the relevance and the popularity of the identified code inefficiency patterns. Our qualitative findings indicate that inefficiencies are diverse and interconnected, affecting multiple aspects of code quality, with logic and performance-related inefficiencies being the most frequent and often co-occurring while impacting overall code quality. Our taxonomy provides a structured basis for evaluating the quality of LLM-generated code and guiding future research to improve code generation efficiency.
@@
LLMs, Code Generation, Code Quality, Python
@@ I Introduction
We conduct our study using HumanEval+  a widely used benchmark for code generation composed of 164 human-crafted Python tasks [6]. We selected three leading open-source LLM models -CodeLlama, DeepSeek-Coder, and CodeGemma- to generate code for the different tasks, resulting in 492 generated code snippets. To evaluate the quality and requirements conformity of the LLM-generated Python code, we employed the higher-capacity GPT-4o-mini model as a judge. The judgments provided initial guidelines for our manual exploratory and qualitative analysis to identify inefficiency patterns in LLM-generated code. We adopted open coding [18] to iteratively derive categories of inefficiencies. Our study results in a taxonomy spanning five categories: General Logic, Performance, Readability, Maintainability, and Errors, comprising 19 subcategories. We validate this taxonomy through a survey with 58 software practitioners and researchers who use LLMs for coding assistance.
@@
  • We study inefficiencies observed in LLM-generated Python code and their prevalence.
  • We propose a taxonomy of inefficiencies in LLM-generated Python code based on their characteristics.
@@ III Methodology
This section presents the methodology used in this study to construct and validate the taxonomy of inefficiencies in LLM-generated Python code. The process consists of five steps, as illustrated in Figure 1. First, we generated code using the HumanEval+ benchmark [6] and collected data from three open-source LLM4Code models: CodeLlama [4], DeepSeek-Coder [21], and CodeGemma [23]. Second, we leveraged a higher-capacity model (GPT-4o-mini) to judge the generated code and identify existing inefficiencies. In the third and fourth steps, we manually analyzed the code samples and the model’s judgments to develop a comprehensive taxonomy. Finally, in the fifth step, we conducted a survey with practitioners and researchers to validate the LLM-generated code inefficiencies identified in our taxonomy. The details of each step are presented in the remainder of this section.
@@ III-A Step 1: Data Selection and Collection
Regarding the data collection, considering the popularity and wide adoption of Python as a programming language [24], this study focused on Python-generated LLM code. Given the focus on Python, we selected HumanEval+ [6] (denoted as $\mathit{A}$ in Figure 1), a widely recognized and extensively employed benchmark for code generation [25, 26, 27, 28]. HumanEval+ enhances the original HumanEval [29] dataset by augmenting it with additional test cases, enabling a more rigorous and reliable assessment of the functional correctness of the generated code [6]. HumanEval+ comprises 146 Python programming tasks, each accompanied by (i) a task description provided as a doc string and (ii) associated unit tests to rigorously evaluate the correctness of the code snippets. Regarding the LLMs, we aimed to select those trained or fine-tuned for code generation tasks to enable a more meaningful and targeted analysis of inefficiencies. To ensure accessibility and feasibility, we established two selection criteria: (i) the models must be open source and (ii) have fewer than 8B parameters. Based on these criteria, we selected the three highest-performing models according to the HumanEval+ leaderboard [30]: CodeLlama-7B [31], CodeGemma-7B [32], and DeepSeek Coder-6.7B [33]. These models, beyond ranking among the top-performing models [30], are also widely adopted for code generation tasks [27]. Although our selection focuses on small open-source models, these models demonstrate performance comparable to proprietary ones, surpassing GPT-3.5 [21] and approaching GPT-4 [34], ensuring representativeness in our study. After defining the setup for our study, we proceeded with generating the code snippets. For the hyperparameters, we adopted a conservative approach by setting the temperature to 0, ensuring deterministic outputs and minimizing randomness in code generation. For prompt construction, we followed the default configuration recommended in the official Hugging Face documentation [32, 31, 33]. This approach ensured a fair assessment of inherent inefficiencies in LLM-generated code rather than artifacts caused by inadequate prompts. As a result of this step, we have a dataset of $164$ generated code snippets for each LLM, resulting in $492$ code snippets ($\mathit{B}$), from the initial dataset of tasks ($\mathit{A}$).
@@ IV-A2 Analysis of Inefficiency Patterns
Sub-Optimal Solution (Memory) 26.83 17.07 12.2 18.70 3.07 3.12
@@
3-Readability 6.71 6.10 1.22 4.67 3.22
Confusing Variable Naming 1.22 1.22 0.61 1.02 2.17 3.00
@@
Syntax Error 0.00 2.44 2.44 1.63 2.29 3.79
Missing Variable Declaration 0.61 0.00 0.61 0.41 2.17 3.60
Missing Module Import 6.10 3.66 1.22 3.66 2.86 3.26
@@
Readability (4.67%) and Errors (5.69%) were less frequent but non-trivial. Syntax Errors were rare ($2.44\%$), confirming the strong syntactic grasp of code LLMs. However, Missing Imports (6.1% in CodeLlama) reveal persistent gaps in dependency resolution, a known limitation even in highly-efficient models like GPT-4 [53].
@@
To gain deeper insights into inefficiency relationships, we analyze how different categories co-occur across models. Figure 3 presents the normalized category co-occurrence heatmap across all models. For each inefficiency category, we calculated the number of times it appeared together with every other category within the same code snippet. These counts were then summed across all models and normalized row-wise by the total occurrences of the corresponding category.
Our results reveal strong inter-dependencies between inefficiency patterns in LLM-generated code. General Logic inefficiencies exhibit the highest co-occurrence rates with other categories, particularly Readability (0.74), Errors (0.64), and Maintainability (0.58). This suggests that flawed logic not only impacts correctness but also contributes to inefficiencies in clarity, structure, and long-term maintainability of the generated code. Similarly, Performance inefficiencies frequently co-occur with Readability (0.48) and Maintainability (0.46), highlighting that performance inefficiencies are often tied to code complexity and structure, rather than isolated algorithmic inefficiencies. Interestingly, Errors demonstrate weaker co-occurrences with other inefficiencies, with their strongest link being to Maintainability (0.36), indicating that while syntax and dependency errors are less common, they may still hinder maintainability efforts by requiring additional debugging and refactoring. Meanwhile, Readability and Maintainability are closely related (0.35), reinforcing the idea that clear code structures contribute to long-term maintainability. These findings suggest that improving LLM-generated logic and performance could lead to broader improvements across multiple inefficiency categories, reducing the need for manual corrections and enhancing code usability in real-world development.
@@ IV-B RQ2: Relevance of Identified Inefficiencies for Practitioners and Researchers
The survey remained open for four weeks, reaching $58$ responses, resulting in a response rate of 7.397.39%, which is in line with rates observed in similar surveys in Software Engineering [39, 54, 55, 56]. First, we present the demographic information about the participants and their experience with LLM-assisted coding. Then, we provide a detailed analysis of the popularity and perceived relevance of inefficiencies in LLM-generated code reported by the participants.
@@ IV-B1 Demographic Information and Experience with Coding using LLMs
All participants answered the demographic questions. We had participants from both academia (60%) and industry (40%). The academic participants consisted of graduate students (29), undergraduate students (three), and postdoctoral researchers/lecturers (three). Industry participants included 11 software developers, seven ML/data science developers, four senior executives, and one cybersecurity developer. About coding experience, 37 participants had more than 5 years of experience, while 11 had between 3-5 years, and 10 had between 1-3 years. LLM-assisted coding is widely adopted, with 98% of the participants using LLMs for coding. Among them, 68% rely exclusively on proprietary models, with ChatGPT and Copilot being the most popular (89.66% and 29.31%, respectively). Regarding open-source, the most popular models are Mistral, the Phi model family, and Qwen-Coder (5.17%, 3.45%, and 3.45%, respectively). Participants reported using a variety of programming languages, with the top five being Python, C/C++, JavaScript, Java, and Go (91.38%, 18.97%, 18.97%, 8.62%, and 3.45%, respectively).
@@
These inefficiencies focus on the challenges/difficulties that users face when interacting with LLMs, rather than the generated code itself. Six participants reported that LLMs struggle with debugging or modifying specific parts of code, a finding consistent with previous studies [57]. In the same way, programming language and library popularity also play a significant role in the efficiency of LLM-generated code. five participants noted that LLMs generate more efficient code for popular languages and libraries like Python and NumPy compared to less common ones like DHL and Tinkter. This difference is likely because LLMs are trained on more examples of popular languages and libraries, making them more familiar with their syntax and usage. Programming language paradigms also play a role—P12111We refer to survey participants using the notation Px, where x is an assigned participant ID. observed that LLMs handle interpreted languages better than compiled ones. Finally, participants expressed varying preferences for chat-based LLMs (e.g., ChatGPT) versus inline code assistants (e.g., Copilot). While P41 found Copilot suggestions as “stupid suggestions and distracted”, P11 preferred inline assistance over chat-based interactions.
@@
Participants also reported various inefficiencies in the generated code that align with the focus of our taxonomy. 19 participants stated that LLMs often struggle to produce the correct logic on the first attempt, even for simple tasks with detailed descriptions, requiring multiple further iterations. P22 specifically noted that LLMs may generate “unnecessarily long or complex code (for simple tasks)” However, P15 found that for more complex tasks such as building a Flask application, LLMs tend to perform “oversimplification to the point of uselessness”, where generation is limited to high-level steps or generalized code that lack specificity, performance efficiency, and functionality. Additionally, three participants reported that handling exceptions and corner cases is often overlooked. These reported inefficiencies fall under the General Logic category of our taxonomy.
@@
Additionally, Six participants highlighted performance issues. Others pointed to a lack of robust, modular, object-oriented architecture and insufficient documentation and comments as well. The inefficiencies reported by the participants align with the literature [16, 14, 16] and fall under Readability, Performance, and Maintainability categories of our taxonomy.
Besides inefficiencies, Seven participants identified context limitations as a key issue. Two participants found that LLMs had a limited context window and frequently forgot past interactions, as mentioned in related work [47]. Meanwhile, P5 reported that LLMs tended to retain outdated context, stating that they needed to “start a new chat to get rid of hallucinations.”
@@ VI Threats to Validity
Internal Validity. A key internal threat to validity is the potential bias in the manual labeling of LLM-generated code. To mitigate such a threat, the first two authors independently labeled the generated code snippets. The inter-rater agreement at the category level is 0.846, which aligns with similar studies in software engineering [39, 13]. Additionally, to further validate our findings, we surveyed 58 participants from both academia and industry, ensuring a broader perspective on the identified inefficiencies. Additionally, we leveraged an LaaJ to assess generated code snippets, providing an additional layer of validation and reducing potential human bias in the labeling process. However, we recognize that LaaJ judgments may have missed certain inefficiencies or introduced biases that could influence human labeling. To mitigate this risk, the two labelers independently reviewed, refined, and expanded upon the LLM’s judgments, ensuring that inefficiencies were identified beyond those initially flagged by the model. Finally, we used GPT-4o-mini to provide initial judgments on code inneficiencies. While all outputs were manually validated to ensure accuracy, the use of such an LLM for initial assessment could have introduced biases or inconsistencies. We mitigated such a threat by reviewing and correcting all code inefficiencies before including them in the final results.
@@ VII Conclusion
VIII Acknowledgment
This work is partially supported by the Fonds de Recherche du Quebec (FRQ), the Canadian Institute for Advanced Research (CIFAR), the Natural Sciences and Engineering Research Council of Canada (NSERC), and the Canada Research Chairs Program. The authors would like to thank all participants who contributed to the survey for their valuable insights. The views and conclusions expressed in this paper are solely those of the authors and do not necessarily reflect those of the supporting organizations.